Evaluate the following indefinite integrals:  


a) ∫ (1/x + 3/x2 - 4/x3 ) dx


b) ∫ (x2 + 2x - 5) / √x dx


c) ∫ x ex dx


1
Expert's answer
2021-05-31T18:18:48-0400

a)


"\int(\dfrac{1}{x}+\dfrac{3}{x^2}-\dfrac{4}{x^3})dx"

"=\ln|x|-\dfrac{3}{x}+\dfrac{2}{x^2}+C"


b)


"\int\dfrac{x^2+2x-5}{\sqrt{x}}dx=\int( x^{3/2}+2x^{1/2}-5x^{-1/2})dx"

"={2 \over 5} x^{5/2}+{2 \over 3} x^{3/2}-10\sqrt{x}+C"


c)


"\int xe^xdx""\int udv=uv-\int vdu"

"u=x, du=dx"


"dv=e^xdx, v=\int e^xdx=e^x"


"\int xe^xdx=xe^x-\int e^xdx"

"=xe^x- e^xdx+C"